A simple random sample of 40 items resulted in a sample mean of 25. The population standard deviation is σ = 5.

Round your answers to two decimal places.

a. What is the standard error of the mean, σx?

b. At 95% confidence, what is the margin of error?
Answer #1

Given

Sample size=n=40

Sample mean=\bar{x}=25

Population Standard deviation=\sigma=5

The the standard error of the mean=

\frac{\sigma}{\sqrt{n}} =\frac{5}{{\sqrt{40}}}=0.7906

The MAriginal error

E=Z_{\alpha /2}\frac{\sigma }{\sqrt{n}}

The critical value at 95% confidencce is 1.96

E=Z_{\alpha /2}\frac{\sigma }{\sqrt{n}}=1.96(0.7906)=1.5495